OpenStudy (anonymous):

solve using the elimination method plz help!!! 5x+3y=-11 7x-2y=19

OpenStudy (anonymous):

Multiply the first equation by 2 and the second equation by 3. Then your y term will be eliminated and you can solve for x. Once you have x, you can plug that value back into either quation and solve for y.

OpenStudy (anonymous):

i get stuck when i get to 5(31/16)+3y=-11

OpenStudy (anonymous):

because i get that x=31/16

OpenStudy (anonymous):

2(5x + 3y = -11) 3(7x - 2y = 19) 10x + 6y = -22 21x - 6y = 57 31x = 35 x = 35/31

OpenStudy (anonymous):

so when i plug it back in it doesnt make sence to me how to solve

OpenStudy (anonymous):

ohhh

OpenStudy (anonymous):

i was messing up with the mulitplicaiton!!!

OpenStudy (anonymous):

5(35/31) + 3y = -11 175/31 + 3y = -11 3y =31(-11) - 175 ------- ---- 31 31 3y = 341 - 175 -------- 3 3y = 166 ---- 3 y = 166 1 --- x --- 3 3 y = 166/9

OpenStudy (anonymous):

but it says the answer is (35/31,-172/31)

OpenStudy (anonymous):

yup y=-172/31 .. just a little typing error in the above solution.

OpenStudy (anonymous):

OOPS.. I'm sorry

OpenStudy (anonymous):

3y = 31(-11) - 175 ------ ---- 31 31 3y = -341 - 175 --------- 31 3y = -516 ---- 31 3y = -516 1 ---- x -- 31 3 y = -516 ---- 93 y = -172 ----- 31